Please provide formulas for the following blank cells.

RETURN RISK CV=RETURN PER RISK
ERus ERpor CHG STDus STDpor CHG CVus CVpor CHG
UK 15% 10% 1.5
SPA 15% 10% 1.5
Answer #1

1)ERpor

Expected Return of portfolio

Erp=Er1*W1+Er2*W2

Erp=15*1/2+15*1/2

=7.5+7.5=15%

CHG change

change UK Nil

change SPA Nil

2)Risk of portfolio

=(SD1)2*W12+(SD2)2*W22+2W1W2(Correlation of stock 1 and 2)*SD1*SD2

3)Return per risk

Sharpe Ratio=Rp-Rf/Risk of portfolio